# A new rotating black hole from the Newman-Janis algorithm

### v1

@@ 3.2 Generalization of the Kerr geometry
    $+\frac{\sin^{2}\theta}{\Sigma(r,\theta)}\left[(r^{2}+a^{2})d\phi-adt\right]^{2}\,,$
    where $\Delta(r)=r^{2}+a^{2}-2Mr+v(r)$ as shown in Ref. [17]. Notice that the metric has almost the same form as the Kerr one except that the function $\Delta(r)$ now includes a correction term, $v(r)$.
  4. 4.
    Finally, we check the Einstein tensor for this geometry (29). The Einstein tensor for this geometry with respect to the basis (26) with $\tilde{A}(r)=0$ takes the diagonal form:
@@
  5. 5.
    Note that the two limits $a\to 0$ and $v(r)\to 0$ are independent. Therefore we can take the $a\to 0$ limit with $v(r)\neq 0$. In this case, the metric (29) provides a new static black hole geometry:
@@
    where $\Delta=r^{2}-2Mr+v(r)$. The source developing this metric satisfies $T_{ab}=G_{ab}/8\pi$ with the Einstein tensor given in Eq. (30). Therefore, we interpret that the metric (29) describes the rotating geometry of this static one.
@@
  6. 6.
    Let us display a few properties of the solution for various special cases.
@@
    • When $v(r)$ is given by Eq. (32), the metric (29) describes the rotating geometry of the static solution (31). The equation of state in this case is $w_{r}=-1$ and $w_{\theta}=w_{\phi}=(w-1)\Sigma(r,\theta)/r^{2}+1$. Note that, from the form of the Einstein tensor (30), the angle dependence of the equation of state is unavoidable unless $v^{\prime\prime}(r)=0$.
    • When $v^{\prime\prime}(r)=v_{c}(v(r)-rv^{\prime}(r))$, i.e.,
@@
      where erf$(x)$ denotes the error function, the equation of state becomes $w_{1}=-1$, and $w_{2}=w_{3}=1+v_{c}\Sigma(r,\theta)/2$. This is an interesting generalization of the Kerr family with a minor modification of the equation of state. Note that $v(r)-rv^{\prime}(r)=v_{2}e^{-v_{c}r^{2}/2}$. Therefore, the stress tensor exponentially decreases with $r$. We analyze the properties of the solution in the next section.
    • Else, the equation of state for the radial part satisfies $w_{1}=-1$. But, $w_{2}=w_{3}$ has an angle dependence. The geometry still describes a stationary rotating black hole with surrounding matter.
As demonstrated in this work, the NJ algorithm does not fully determine the geometry of the system, but rather leaves room for imposing physical requirements. It determines the geometry up to Eq. (25). The final form of the metric is determined by imposing the physical condition (the axially rotating condition $R_{\hat{r}\hat{\theta}}=0$) and selecting an appropriate equation of state in the last two steps.
3.3 The inconsistent choices
In this subsection, we demonstrate that arbitrary choices fail to yield a physically consistent geometry by illustrating two examples of metric functions that do not align with axially symmetric rotation in Einstein’s theory.
  1. 1.
    Consider the choice $q(x)=x$, $p=0$. $\Sigma(r,\theta)=r^{2}+a^{2}\cos^{2}\theta$. The metric with this choice becomes,
    $ds^{2}=-\frac{\Sigma(r,\theta)\Delta(r)}{\Gamma^{2}(r)}dt^{2}+\Sigma(r,\theta)d\theta^{2}+\frac{\Sigma(r,\theta)}{\Delta(r)}dr^{2}+\Sigma(r,\theta)\sin^{2}\theta\left(d\phi+\frac{a}{\Gamma(r)}dt\right)^{2}\,.$ (34)
    When we calculate the Ricci tensor, we can easily find that there is no way to set $R_{\hat{r}\hat{\theta}}=-\frac{6a^{2}r\sin\theta\cos\theta\sqrt{\Delta(r)}}{\Sigma^{3}(r,\theta)}$ to vanish for all $\theta$ unless $a\neq 0$.
  2. 2.
    We next consider the choice $q(x)=x$ and leave $p(x)$ free. Therefore, the metric becomes
    $ds^{2}$ $=$ $-\frac{\Sigma(r,\theta)\Delta(r)}{\left(\Gamma(r)-{\dot{p}}(x)a^{2}\sin^{2}\theta\right)^{2}}\left(dt-{\dot{p}}(x)a\sin^{2}\theta d\phi\right)^{2}+\Sigma(r,\theta)d\theta^{2}+\frac{\Sigma(r,\theta)}{\Delta(r)}dr^{2}$ (35)
    $+\frac{\Sigma(r,\theta)\sin^{2}\theta}{\left(\Gamma(r)-{\dot{p}}(x)a^{2}\sin^{2}\theta\right)^{2}}(\Gamma(r)d\phi-adt)^{2}\,.$
    Here, $\Delta(r)=r^{2}+a^{2}-2Mr+v(r)$ and $\Gamma(r)\equiv r^{2}+a^{2}+\tilde{A}(r)$.
    The Ricci component $R_{r\theta}$ becomes
    $R_{\hat{r}\hat{\theta}}$ $=$ $\frac{3a^{2}\sqrt{\Delta(r)}r\sin 2\theta}{2\rho^{6}\left(\Gamma(r)-a^{2}{\dot{p}}(x)\sin^{2}\theta\right)^{2}}\Big{[}4a^{2}\tilde{A}(r){\dot{p}}(x)\sin^{2}\theta+2(a^{2}+r^{2})^{2}-2(a^{2}+r^{2}+\tilde{A}(r))^{2}$ (36)
    $+2({\dot{p}}(x)-1)\left(\left(a^{2}+r^{2}\right)^{2}-p^{\prime}(x)a^{4}\sin^{4}\theta\right)$
    $\left.+4\rho^{4}\frac{-{\ddot{p}}(x)\left(\tilde{A}^{\prime}(r)+2r\right)\cos\theta+2{\dot{p}}(x)\tilde{A}^{\prime}(r)}{8r}\right]\,.$
    It appears that a unique way which makes $R_{\hat{r}\hat{\theta}}=0$ for all $r$ and $\theta$ must be $\tilde{A}(r)=0$ and ${\dot{p}}(x)=1$. This result reduces the metric into the generalized Kerr solution form (29).
@@ 4 Analysis of the new solution
$\rho^{s}=\frac{v_{2}e^{-v_{c}r^{2}/2}}{r^{4}}\,,~{}~{}p^{s}_{\hat{r}}=-\rho^{s}\,,~{}~{}p^{s}_{\hat{\theta}}=p^{s}_{\hat{\phi}}=\left(1+\frac{v_{c}r^{2}}{2}\right)\rho^{s}\,.$ (41)
@@ 6 Summary and discussions
We investigated the Newman-Janis (NJ) algorithm in a general context as a method for constructing rotating geometries from static ones. Our analysis demonstrated that the NJ algorithm serves as a powerful mathematical tool, simplifying the structure of general rotating metrics by imposing basic physical constraints. Through a detailed examination of the black hole metric to which the NJ algorithm is applied, we derived a new rotating black hole solution in a quite general form.
While the traditional approach applies the NJ algorithm to a static, spherically symmetric black hole to obtain a rotating counterpart, our method first derives the rotating black hole geometry directly through an analysis of the NJ algorithm. We then work backward to retrieve the corresponding spherically symmetric black hole solution. One key requirement for a well-defined rotating black hole solution is that the Einstein tensor component $R_{\hat{r}\hat{\theta}}$ should vanish in an orthonormal basis, a condition that our newly derived metric satisfies. The NJ algorithm is particularly well-defined when applied to static, spherically symmetric black holes with $f(r)=g(r)$.
@@
Notably, we found that the geometry of the rotating black hole solution is not uniquely determined by asymptotic parameters such as mass, charge, angular momentum, or any other gauge charges. Specifically, a matter field exists near the black hole horizon, with its density decaying not later than exponentially at large distances. This solution, therefore, challenges the no-hair conjecture and suggests the existence of a new form of black hole ”hair” that remains undetectable in the asymptotic region. Investigating the properties of such a matter field in curved spacetime would be an intriguing direction for future research. An alternative, indirect approach is to analyze black hole thermodynamics by defining a new charge associated with this field and identifying the corresponding potential.
@@ Acknowledgments
We are grateful to Wontae Kim and Stefano Scopel for their hospitality during our visit to the Workshop on Cosmology and Quantum Spacetime (CQUeST 2024), and Inyong Cho to the workshop on How to use AI in Astrophysics Theory. H.-C. Kim (RS-2023-00208047) and W. Lee (RS-2022-NR075087, CQUeST: RS-2020-NR049598) were supported by Basic Science Research Program through the National Research Foundation of Korea funded by the Ministry of Education.

### v2

@@ 2 Rotating geometries from the static spherically symmetric black hole
$+\frac{\sin^{2}\theta}{\Sigma(r,\theta)}\left[adt-(r^{2}+a^{2})d\phi\right]^{2}\,,$
where $\Delta(r)=r^{2}+a^{2}-2Mr+v(r)$ as shown in Refs. [75, 18, 21]. In Ref. [75], the Kerr-Schild coordinate was analyzed in detail. Notice that the metric has almost the same form as the Kerr one except that the function $\Delta(r)$ now includes a correction term, $v(r)$.
Still, we cannot say that the metric (10) to describe a solution of the Einstein equation. To ensure the metric to be a solution, one should prescribe what matter make the geometry. For this purpose, we check the Einstein tensor and check the equation of state required to specify a solution. We take the tetrad frame [72, 73, 21]
$e^{\mu}_{\hat{t}}$ $=$ $\frac{(r^{2}+a^{2},0,0,a)}{\sqrt{\Sigma\triangle}}\,,~{}~{}~{}~{}e^{\mu}_{\hat{r}}=\frac{\sqrt{\triangle}(0,1,0,0)}{\sqrt{\Sigma}}\,,$
$e^{\mu}_{\hat{\theta}}$ $=$ $\frac{(0,0,1,0)}{\sqrt{\Sigma}}\,,\quad e^{\mu}_{\hat{\phi}}=-\frac{(a\sin^{2}\theta,0,0,1)}{\sqrt{\Sigma}\sin\theta}\,.$ (11)
The Einstein tensor for this geometry with respect to the basis (2) with $\tilde{A}(r)=0$ takes the diagonal form:
@@
Note that the two limits $a\to 0$ and $v(r)\to 0$ are independent. Therefore we can take the $a\to 0$ limit with $v(r)\neq 0$. In this case, the metric (10) provides a new static geometry:
@@
where $\Delta=r^{2}-2Mr+v(r)$. The source developing this metric satisfies $T_{\mu\nu}=G_{\mu\nu}/8\pi$ with the Einstein tensor given in Eq. (12). Therefore, we interpret that the metric (10) describes the rotating geometry of this static one.
@@
Let us display a few properties of the solution (10) for various special cases.
@@
  • When $v(r)$ is given by Eq. (14), the metric (10) describes the rotating geometry of the static solution (13). The equation of state in this case is $w_{r}=-1$ and $w_{\theta}=w_{\phi}=(w-1)\Sigma(r,\theta)/r^{2}+1$. When $w=1$, the solution correspond to the Kerr-Newman solution. Note that, from the form of the Einstein tensor (12), the angle dependence of the equation of state is unavoidable unless except for the above three cases.
  • Any functional dependence for $v$ can be constructed by imposing the stress tensor correspondingly once it satisfies the physically acceptable conditions such as asymptotic flatness. However, to accept as an interesting generalization of Kerr family, the geometry should be constructed from a minor modification of the equation of state for matter. A simplest unnotices equation of state $w_{1}=-1$, and $w_{2}=w_{3}=1+v_{c}\Sigma(r,\theta)/2$ is achieved when $v^{\prime\prime}(r)=v_{c}(v(r)-rv^{\prime}(r))$. In this case,
@@
    where erf$(x)$ denotes the error function. This is an interesting generalization of the Kerr family with a minor modification of the equation of state. Note that $v(r)-rv^{\prime}(r)=v_{2}e^{-v_{c}r^{2}/2}$. Therefore, the stress tensor exponentially decreases with $r$. We analyze the properties of the solution in the next section.
  • Else, the equation of state for the radial part satisfies $w_{1}=-1$. But, $w_{2}=w_{3}$ has an angle dependence. The geometry still describes a stationary rotating black hole with surrounding matter. An interesting observation here is that because the stress tensor is linear in $v$, any combinations of the above cases also belongs to a solution of the Einstein equation. For example, one may add $v_{1}=q^{2}$ [electric field], $v_{2}=Kr^{-2}$ [$w=2$ case in Eq. (14)], and $v_{3}$ [the above erf functions].
Starting from the metric ansatz (10), we imposes the axially rotating condition $R_{\hat{r}\hat{\theta}}=0$ and select an appropriate equation of state to determine the final form of the metric.
@@ 3 Analysis of the new solution
$\rho^{s}=\frac{v_{2}e^{-v_{c}r^{2}/2}}{8\pi r^{4}}\,,~{}~{}p^{s}_{\hat{r}}=-\rho^{s}\,,~{}~{}p^{s}_{\hat{\theta}}=p^{s}_{\hat{\phi}}=\left(1+\frac{v_{c}r^{2}}{2}\right)\rho^{s}\,.$ (20)
The energy density of the new matter field is similar to that of an electric field, $\rho_{c}=\frac{Q^{2}}{8\pi r^{4}}$, but it has the property that its numerator decays exponentially as a function of the square of $r$, rather than as a constant. The radial pressure equals the negative energy density, resulting in a matter field that can coexist with the black hole outside the event horizon. The transverse pressures have an additional contribution from the presence of $v_{c}$.
@@ 5 Summary and discussions
We investigated the Newman-Janis (NJ) algorithm [10] and used results of the “non-complexification procedure” [13] as a method for constructing rotating geometries from static ones. To describe the geometry of a rotating black hole, we solved the Einstein equations, starting with Eq. (2). We first demand the Einstein tensor component $R_{\hat{r}\hat{\theta}}$ vanishes so that the geometry is consistent with the axial symmetry. We then, check the Einstein tensor to get proper equation of states the stress tensor should follow.
@@
Notably, we found that the geometry of the rotating black hole solution is not uniquely determined by asymptotic parameters such as mass, charge, angular momentum, or any other gauge charges. Specifically, a matter field exists near the black hole horizon, with its density decaying not later than exponentially at large distances. This solution, therefore, challenges the no-hair conjecture and suggests the existence of a new form of black hole“hair” that remains undetectable in the asymptotic region. Investigating the properties of such a matter field in curved spacetime would be an intriguing direction for future research. An alternative, indirect approach is to analyze black hole thermodynamics by defining a new charge associated with this field and identifying the corresponding potential.
@@ Acknowledgments
We are grateful to Wontae Kim and Stefano Scopel for their hospitality during our visit to the Workshop on Cosmology and Quantum Spacetime (CQUeST 2024), and Inyong Cho to the workshop on How to use AI in Astrophysics Theory. H.-C. Kim (RS-2023-00208047) and W. Lee (RS-2022-NR075087, CQUeST: RS-2020-NR049598) were supported by Basic Science Research Program through the National Research Foundation of Korea funded by the Ministry of Education. We thank Metin Gurses for giving valuable comments.